OpenStudy (anonymous):

Factor completely 5w^3-20w^2+3w-12

OpenStudy (anonymous):

you could try factor by grouping which means you would factor 5w^3 and -20w^2 together. and 3w and -12 together

OpenStudy (anonymous):

so it would be...-12w-12?? o my im lost

OpenStudy (anonymous):

well lets take 5w^3 -20w^2 for now. what do the numbers 5 and 20 have in common? like you have to see if they can both be 'factored' by the same number.

OpenStudy (anonymous):

5*4=20

OpenStudy (anonymous):

yeah that's right! so we can 'take out 5 in common' or factor the 5. basically factoring is just a simpler way of writing the same equation. so 5w^3 -20w^2 = 5(w^3 - 4w^2)

OpenStudy (anonymous):

because brackets mean multiplication right? so if you multiplied the 5 you would get the same thing

OpenStudy (mathstudent55):

Factor by grouping. See what you can factor out of the first two terms, and see what you can factor out of the last two terms. 5w^3 - 20w^2 + 3w - 12 first two terms have 5w^2 as common factor last two terms have factor of 3 5w^2(w - 4) + 3(w - 4) Now notice that (w - 4) is a common factor, so take it out (w - 4)(5w^2 + 3)

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Yes, w - 4 cannot be factored, and 5w^2 + 3 cannot be factored, so that's it.